OpenStudy (anonymous):

The surface area of a rectangular prism is 15 m2. What is the surface area of this prism, measured in cm2?

OpenStudy (anonymous):

@TheOcean

OpenStudy (anonymous):

@hippo4

OpenStudy (anonymous):

@asatt32

OpenStudy (anonymous):

@ruthym

OpenStudy (anonymous):

@Pamela16

OpenStudy (anonymous):

@Loser66 @Kainui @SeaTurtle113 @cowgirlcrazy66 @wolf1728 @AravindG @vshiroky @zepdrix @nincompoop @xoxo_nanci @xMiss_Insanityx

OpenStudy (anonymous):

A. 0.0015 cm2 B. 150,000 cm2 C. 1500 cm2 D. 0.15 cm2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

well i needhelp on this cuz idk how to do this and idk if you have done this yet

OpenStudy (anonymous):

yea i did can u help me with mine after

OpenStudy (loser66):

first off, convert m^2 to cm^2 100cm = 1 m, ok? so, (100cm)^2 = 1m^2 ok?

OpenStudy (anonymous):

okay

OpenStudy (loser66):

you have the ratio: \[\dfrac{10000cm^2}{1m^2}\]

OpenStudy (loser66):

so, just time 15m^2 to get 150000cm^2 . Dat sit